### v1

@@ Abstract
Swarm UAV autonomous flight for Long-Horizon (LH) tasks is crucial for advancing the low-altitude economy. However, existing methods focus only on specific basic tasks due to dataset limitations, failing in real-world deployment for LH tasks. LH tasks are not mere concatenations of basic tasks, requiring handling long-term dependencies, maintaining persistent states, and adapting to dynamic goal shifts. This paper presents U2UData-2, the first large-scale swarm UAV autonomous flight dataset for LH tasks and the first scalable swarm UAV data online collection and algorithm closed-loop verification platform. The dataset is captured by 15 UAVs in autonomous collaborative flights for LH tasks, comprising 12 scenes, 720 traces, 120 hours, 600 seconds per trajectory, 4.32M LiDAR frames, and 12.96M RGB frames. This dataset also includes brightness, temperature, humidity, smoke, and airflow values covering all flight routes. The platform supports the customization of simulators, UAVs, sensors, flight algorithms, formation modes, and LH tasks. Through a visual control window, this platform allows users to collect customized datasets through one-click deployment online and to verify algorithms by closed-loop simulation. U2UData-2 also introduces an LH task for wildlife conservation and provides comprehensive benchmarks with 9 SOTA models. U2UData-2 can be found at https://fengtt42.github.io/U2UData-2/.
1 Introduction
@@
Figure 1: U2UData-2 collects a large-scale swarm UAV autonomous flight dataset for LH tasks. U2UData-2 also bulid a scalable swarm UAV data online collection and algorithm closed-loop verification platform, supporting the customization of simulators, UAVs, sensors, flight algorithms, formation modes, and LH tasks. Through a visual control window, U2UData-2 allows users to collect customized datasets through one-click deployment online and to verify algorithms by closed-loop simulation.
@@
Swarm Unmanned Aerial Vehicle (UAV) autonomous flight Wang et al. (2020) can solve the inherent limitations of single-UAV through collaborative perception, localization, communication, navigation, tracking, and task re-allocation. By leveraging UAV-to-UAV (U2U) technologies, swarm UAVs overcome single-viewpoint occlusion and sensor range limits through multi-view collaborative perception Xu et al. (2022b). Furthermore, swarm UAV ensures operational robustness against failures or obstacles for accurate navigation Li et al. (2021) and dynamic tracking Liu et al. (2020) by collaborative localization, communication, and task re-allocation, while mitigating computational constraints via shared processing and decentralized decision-making. Finally, swarm UAV autonomous flight can achieve robust, scalable, and adaptive task execution in complex and harsh environments unattainable by single-UAV systems.
@@
Swarm UAV autonomous flight for Long-Horizon (LH) tasks is crucial for advancing the low-altitude economy. LH tasks Feng et al. (2025b) are complex, multi-step tasks that require sustained planning, sequential decision-making, and extended execution over a prolonged period to achieve a final goal. The practical applications of swarm UAV are almost all LH tasks, such as logistics distribution Betti Sorbelli (2024), patrol security Yuan et al. (2024), wildlife conservation Feng et al. (2024b), disaster rescue Sun et al. (2024), and infrastructure inspection Pan et al. (2024).
@@
However, existing methods focus only on specific basic tasks due to dataset limitations, failing in real-world deployment for LH tasks. Existing swarm UAV flight datasets, as shown in Table 1, CoPerception-UAVs Hu et al. (2022) and CoPerception-UAVs+ Hu et al. (2023) are based on open-source simulators such as AirSim Shah et al. (2018) and CARLA Dosovitskiy et al. (2017) and consider only 1 terrain, 1 weather, and 1 to 2 sensor types; they collect datasets using fixed altitude and consistent or fixed formation mode. In real-world scenarios, compared to autonomous driving Liu et al. (2024), autonomous flight has more freedom, faces more complex environments, and is more susceptible to the influence of temperature, humidity, and airflow due to its smaller size. Obviously, there will be a clear domain gap between existing synthetic data and real-world data. U2UData Feng et al. (2024b) is the first swarm UAVs autonomous flight dataset, which is collected by 3 UAVs flying autonomously in the U2USim Han et al. (2024), covering 4 terrains, 7 weather conditions, and 8 sensor types. Due to the emergence of U2UData, swarm UAV autonomous flight algorithms have begun to be studied. But U2UData is hard to use for exploring LH tasks: 1) the length of each trajectory in U2UData is only 15 seconds and only focuses on basic collaborative perception and tracking tasks; 2) the dataset size, tasks, and settings are preset and fixed and cannot be expanded. LH tasks are not mere concatenations of basic tasks, requiring handling long-term dependencies, maintaining persistent states, and adapting to dynamic goal shifts. Therefore, building a scalable swarm UAV autonomous flight dataset for LH tasks is an urgent and challenging work for real-world deployment.
@@
In this paper, we present U2UData-2, as shown in Figure 1, the first large-scale swarm UAV autonomous flight dataset for LH tasks and the first scalable swarm UAV data online collection and algorithm closed-loop verification platform. 1) The dataset is captured by 15 UAVs in autonomous collaborative flight for LH tasks (dataset size: 3.62T), comprising 12 scenes (weather and terrain combination), 720 traces, 120 hours (each trace 600 seconds), 4.32M LiDAR frames, 12.96M RGB, and 12.96M depth frames. This dataset also includes brightness, temperature, humidity, smoke, and airflow values covering all flight routes. 2) The platform supports the customization of simulators, UAVs, sensors, flight algorithms, formation modes, and LH tasks. Through a visual control window, this platform allows users to collect customized datasets through one-click deployment online and to verify algorithms by closed-loop simulation, which can greatly alleviate the limitations of existing datasets on algorithm development. 3) U2UData-2 also introduces an LH task for wildlife conservation and provides 9 state-of-the-art swarm algorithms for benchmarking. All datasets, platforms, benchmarks, and video tutorials have been open-sourced and are available for public use. Our contributions can be summarized as follows:
@@
  • Dataset [^1] https://huggingface.co/datasets/fengtt42/U2UData-2/tree/main. We collect the first swarm UAV autonomous flight dataset for LH tasks, with a size of over 3.62T.
  • Platform [^2] https://huggingface.co/datasets/fengtt42/U2USim-2/tree/main. We build the first scalable swarm UAV data online collection and algorithm closed-loop verification platform, which allows users to collect customized datasets through one-click deployment online and to verify algorithms by closed-loop simulation.
  • Benchmark [^3] https://huggingface.co/fengtt42/U2UData-2-Client/tree/main. We introduce an LH task for wildlife conservation and provide comprehensive benchmarks with 9 SOTA models.
2 Related Work
@@
Swarm UAV autonomous flight. Current low-altitude economy research mainly focuses on single-UAV autonomous flight and has matured core capabilities Zhang et al. (2022); Xu et al. (2022a); Feng et al. (2024a), including object detection, semantic segmentation, localization, obstacle avoidance, navigation, tracking, and stabilized flight control in controlled environments. However, they still suffer from many real-world challenges, for example: (1) Their perception remains fundamentally constrained by single-viewpoint occlusion and limited sensor range Hu et al. (2023); Lu et al. (2024), severely reducing situational awareness in dynamic open environments. (2) Onboard computational resources restrict real-time decision-making for dynamic obstacle negotiation and executing LH tasks Feng et al. (2025a); Valente et al. (2024); Shen et al. (2025). (3) Operational robustness is inherently fragile Sun et al. (2023); Zhao et al. (2023), as hardware failures or unexpected obstacles often lead to task failure with no redundancy. Swarm UAV autonomous flight Wang et al. (2020) can solve the inherent limitations of single-UAV through collaborative perception, localization, communication, navigation, tracking, and task re-allocation. Due to the lack of datasets, research on autonomous flight algorithms for swarm UAV has just begun.
@@
Swarm UAV simulators. Existing swarm UAV simulators include FightGear [^4] https://www.flightgear.org/, XPlan [^5] https://www.x-plane.com/, Jmavsim [^6] https://docs.px4.io/main/en/sim_jmavsim/, Gazebo [^7] https://gazebosim.org/home, AirSim Shah et al. (2018), Rfly-Sim [^8] https://rflysim.com/doc/zh/, Isaac Sim [^9] https://developer.nvidia.com/isaac-sim, and U2USim Han et al. (2024). Swarm UAV simulators need to more realistically simulate dynamic physical characteristics (such as collision); sensors such as IMU, camera, GPS, LiDAR, temperature, humidity, and airflow due to their small size; and interaction with the ROS ecosystem. FightGear is not open source. XPlan and Jmavsim can only interact with ROS. Gazebo, AirSim, and RflySim can interact with ROS, simulate physical collision, and output visual sensor content. AirSim and RflySim can also implement weather control. However the information on these simulators is purely simulated, and the models trained on these simulators are difficult to run in the real world. Isaac Sim and U2USim add real environment data based on previous simulators. Isaac Sim can visually realize digital twins of the real world through GPU rendering, but it is difficult to provide modal information other than visual and LiDAR modalities. U2USim is the first real-world mapping swarm UAV simulator, taking Yunnan Province as the prototype, including 4 terrains, 7 weather conditions, and 8 sensor types. However, all parameters of U2USim are fixed: it only contains 3 types of animals, the number of animals is fixed, the intensity and range of weather are fixed, and the take-off point of the UAV is also fixed. If we want to test in another terrain, we need to fly to the target location for a long time before each test.
Figure 2: Swarm animal behavior characterization. This platform can represent complex swarm dynamics and ecological relationships between animals. It accurately models interactions such as predator-prey dynamics and competition for resources, showing how these relationships influence collective movement patterns.
@@
Swarm UAV datasets. Public swarm UAV datasets have significantly accelerated progress in UAV flight technologies in recent years. As shown in Table 1, existing swarm UAV datasets include CoPerception-UAVs Hu et al. (2022), CoPerception-UAVs+ Hu et al. (2023), and U2UData Feng et al. (2024b). CoPerception-UAVs Hu et al. (2022) and CoPerception-UAVs+ Hu et al. (2023) rely on open-source simulators like AirSim Shah et al. (2018) and CARLA Dosovitskiy et al. (2017), featuring limited terrain, weather, and sensor types. These datasets collect data at fixed altitudes and in consistent or fixed formation modes. In contrast to autonomous driving Liu et al. (2024), UAVs’ autonomous flight presents greater freedom, encounters more complex environments, and is more susceptible to the influence of temperature, humidity, and airflow due to its smaller size. Hence, there exists a notable domain gap between existing synthetic data and real-world data, potentially limiting the generalization of models trained. U2UData Feng et al. (2024b) is the first large-scale cooperative perception dataset for swarm UAVs autonomous flight, which is collected by three UAVs flying autonomously in the U2USim Han et al. (2024), covering a 9 km2 flight area, 4 terrains, 7 weather conditions, and 8 sensor types. U2UData manually selects 100 scenarios for each weather condition; U2UData collects 15 seconds of swarm UAV cooperative perception dataset for each scenario. Due to the emergence of U2UData, swarm UAV autonomous flight algorithms have begun to be studied. However, since U2UData only considers three UAVs tracking three animals, the length of each trajectory is only 15 seconds, and the dataset size and setting is fixed and cannot be expanded; only basic collaborative perception and tracking tasks can be designed. Complex LH tasks Liu et al. (2025) for swarm UAVs in dynamic open environments cannot be explored.
3 U2UData-2 Platform
U2UData-2 bulids the first scalable swarm UAV data online collection and algorithm closed-loop verification platform, as shown in Figure 1, supporting the customization of simulators, UAVs, sensors, flight algorithms, formation modes, and LH tasks. Through a visual control window, U2UData-2 allows users to collect customized datasets through one-click deployment online and to verify algorithms by closed-loop simulation. We have built a video tutorial for this platform. For each part of the scalable operation, users can complete it one by one according to the video tutorial.
Real-world mapping simulator. The platform is based on U2USim Han et al. (2024), a real-world mapping swarm UAV simulator.The platform uses Unreal Engine (UE) 5.2101010https://www.unrealengine.com/en-US/unreal-engine-5 to construct a scaled-down 3km*3km simulated environment map based on the map of Yunnan Province. The platform includes 4 types of terrain: mountains, hills, plains and basins. The elevation range is [56.6, 3000]m. Based on the vegetation and animal distribution in Yunnan, 58 types of original forest vegetation and 15 types of animal assets were constructed, and more than 15 superposition methods were used to combine vegetation assets, including epiphytic growth, diagonal staggered growth, and so on. Among them, the leaves of each plant will dynamically change with wind, snow, and other weather conditions. As shown in Figure 2, this platform can represent complex swarm dynamics and ecological relationships between animals. It accurately models interactions such as predator-prey dynamics and competition for resources, showing how these relationships influence collective movement patterns. The platform includes 7 weather conditions: sunny, rain, snow, sandstorm, wind, thunder, and fog at specific positions within the simulation environment. The platform uses the real meteorological data of Yunnan Province collected by the China Meteorological Center to map the simulation environment based on longitude and latitude. Among them, temperature and humidity are scalars, and missing values are filled by the moving average method (interval 5m). Wind speed and wind direction are first decomposed into scalars along longitude and latitude, then missing values are filled by sliding average, and finally constructed by vector synthesis.
@@ 3 U2UData-2 Platform
Figure 3: The scalable swarm UAV data online collection and algorithm closed-loop verification platform. Wind throughout the map.
@@
Scalable LH tasks. The platform provides four preset LH tasks: wildlife conservation employs adaptive animal tracking algorithms using real-time behavior prediction across variable terrains and vegetation density. Logistics distribution dynamically reroutes paths around simulated urban obstacles and weather disruptions while maintaining payload integrity. Patrol security implements anomaly detection through continuous environmental scanning, adapting surveillance patterns to emergent threats in real-time. Disaster rescue prioritizes survivor identification in volatile conditions (collapsing structures, spreading fires) via multi-sensor fusion and probabilistic hazard mapping. Each task integrates specialized perception-action loops that respond to unpredictable environmental changes without predefined waypoints, such as sudden weather shifts or moving obstacles. New LH tasks (e.g., precision agriculture) can be added by modifying the UE5.2 simulator source code.
Scalable flight algorithms and formation modes. The platform supports four swarm UAV autonomous flight algorithms for four LH tasks: wildlife conservation, logistics distribution, patrol security, and disaster rescue. Those algorithms are built upon modularized components, including task planning, collaborative perception, localization, communication, navigation, tracking, and task re-allocation. Users can add or modify these algorithms via the open-source code of the visual control window, where modular code blocks allow drag-and-drop replacement or augmentation of existing logic. New autonomous flight algorithms for custom LH tasks can be integrated by directly modifying the provided Python/ROS 2 interfaces in the code repository. The platform implements three distinct swarm formation modes: Discipline formation mode maintains strict geometric coordination (e.g., linear/radial arrays) for high-precision collaborative tasks, with real-time position correction compensating for environmental disturbances. Fixed formation mode enables individual UAVs to follow predefined paths, critical for infrastructure inspection or convoy protection scenarios. Autonomous formation mode supports dynamic reconfiguration where UAVs independently adapt spacing and topology using real-time perception data, ideal for complex environments like wildlife conservation or disaster rescue. Users can select any swarm formation modes via the visual control window for task-specific optimization.
@@
Table 2: A detailed comparison of the data size between U2UData-2 with existing swarm UAV datasets.
@@
U2UData-2 12.96M 1920*1080 12.96M 4.32M 25.92M 12.96M 12.96M 12.96M 12.96M
Visual control window. The platform provides a visual control window, as shown in Figure 4. Users can collect datasets through a one-click deployment of the customized UAV starting point, weather, LH tasks, swarm UAV autonomous flight algorithms, and swarm formation mode. Users can also verify swarm UAV autonomous flight algorithms by closed-loop simulation. For the platform basic capability test, users can first click the ”keyboardControl” button, and then control the UAV by clicking the following buttons: ”Take off”, ”Land”, ”Up”, ”Down”, ”Move Forward”, ”Move Backward”, ”Move Left”, and ”Move Right”. We’ve also implemented XBOX controller control. Connect your XBOX and open the simulator to directly control the UAV with the controller. It’s important to note that XBOX controller control and keyboard control are mutually exclusive.
4 U2UData-2 Dataset
U2UData-2 collects the first swarm UAV autonomous flight dataset for LH tasks, with a size of over 3.62T. The dataset is collected by 15 UAVs in autonomous formation mode for the LH task (wildlife conservation).
LH tasks. U2UData-2 only collects one LH task: wildlife conservation, with a size of over 3.62T. Due to the huge amount of data, users can collect datasets for other LH tasks on the U2UData-2 platform on their own.
@@ 4 U2UData-2 Dataset
Table 4: Data collection settings between U2UData-2 with existing swarm UAV datasets. ET-Length: The length of each trajectory. TNT: The total length of trajectories.

### v2

@@ Abstract
Swarm UAV autonomous flight for Embodied Long-Horizon (ELH) tasks is crucial for advancing the low-altitude economy. However, existing methods focus only on specific basic tasks due to dataset limitations, failing in real-world deployment for ELH tasks. ELH tasks are not mere concatenations of basic tasks, requiring handling long-term dependencies, maintaining embodied persistent states, and adapting to dynamic goal shifts. This paper presents U2UData+, the first large-scale swarm UAV autonomous flight dataset for ELH tasks and the first scalable swarm UAV data online collection and algorithm closed-loop verification platform. The dataset is captured by 15 UAVs in autonomous collaborative flights for ELH tasks, comprising 12 scenes, 720 traces, 120 hours, 600 seconds per trajectory, 4.32M LiDAR frames, and 12.96M RGB frames. This dataset also includes brightness, temperature, humidity, smoke, and airflow values covering all flight routes. The platform supports the customization of simulators, UAVs, sensors, flight algorithms, formation modes, and ELH tasks. Through a visual control window, this platform allows users to collect customized datasets through one-click deployment online and to verify algorithms by closed-loop simulation. U2UData+ also introduces an ELH task for wildlife conservation and provides comprehensive benchmarks with 9 SOTA models.
Datasethttps://fengtt42.github.io/U2UData-2/
Introduction
@@
Figure 1: U2UData+ collects a large-scale swarm UAV autonomous flight dataset for ELH tasks. U2UData+ also bulid a scalable swarm UAV data online collection and algorithm closed-loop verification platform, supporting the customization of simulators, UAVs, sensors, flight algorithms, formation modes, and ELH tasks. Through a visual control window, U2UData+ allows users to collect customized datasets through one-click deployment online and to verify algorithms by closed-loop simulation.
@@
Swarm Unmanned Aerial Vehicle (UAV) autonomous flight (12) can solve the inherent limitations of single-UAV through collaborative perception, localization, communication, navigation, tracking, and task re-allocation. By leveraging UAV-to-UAV (U2U) technologies, swarm UAVs overcome single-viewpoint occlusion and sensor range limits through multi-view collaborative perception (14). Furthermore, swarm UAV ensures operational robustness against failures or obstacles for accurate navigation (15) and dynamic tracking (16) by collaborative localization, communication, and task re-allocation, while mitigating computational constraints via shared processing and decentralized decision-making. Finally, swarm UAV autonomous flight can achieve robust, scalable, and adaptive task execution in complex and harsh environments unattainable by single-UAV systems.
@@
Swarm UAV autonomous flight for Embodied Long-Horizon (ELH) tasks is crucial for advancing the low-altitude economy. ELH tasks (feng2025evoagent) are complex, multi-step tasks that require sustained embodied planning, sequential decision-making, and extended execution over a prolonged period to achieve a final goal. The practical applications of swarm UAV are almost all ELH tasks, such as logistics distribution (betti2024uav), wildlife conservation (feng2024u2udata), disaster rescue (sun2024joint), and infrastructure inspection (pan2024unmanned).
@@
However, existing methods focus only on specific basic tasks due to dataset limitations, failing in real-world deployment for ELH tasks. Existing swarm UAV flight datasets, as shown in Table 1, CoPerception-UAVs (17) and CoPerception-UAVs+ (36) are based on open-source simulators such as AirSim (airsim) and CARLA (CARLA) and consider only 1 terrain, 1 weather, and 1 to 2 sensor types; they collect datasets using fixed altitude and consistent or fixed formation mode. In real-world scenarios, compared to autonomous driving (liu2024survey), autonomous flight has more freedom, faces more complex environments, and is more susceptible to the influence of temperature, humidity, and airflow due to its smaller size. Obviously, there will be a clear domain gap between existing synthetic data and real-world data. U2UData (feng2024u2udata) is the first swarm UAVs autonomous flight dataset, which is collected by 3 UAVs flying autonomously in the U2USim (u2usim), covering 4 terrains, 7 weather conditions, and 8 sensor types. Due to the emergence of U2UData, swarm UAV autonomous flight algorithms have begun to be studied. But U2UData is hard to use for exploring ELH tasks: 1) the length of each trajectory in U2UData is only 15 seconds and only focuses on basic collaborative perception and tracking tasks; 2) the dataset size, tasks, and settings are preset and fixed and cannot be expanded. ELH tasks are not mere concatenations of basic tasks, requiring handling long-term dependencies, maintaining persistent states, and adapting to dynamic goal shifts. Therefore, building a scalable swarm UAV autonomous flight dataset for ELH tasks is an urgent and challenging work for real-world deployment.
@@
In this paper, we present U2UData+, as shown in Figure 1, the first large-scale swarm UAV autonomous flight dataset for ELH tasks and the first scalable swarm UAV data online collection and algorithm closed-loop verification platform. 1) The dataset is captured by 15 UAVs in autonomous collaborative flight for ELH tasks (dataset size: 3.62T), comprising 12 scenes (weather and terrain combination), 720 traces, 120 hours (each trace 600 seconds), 4.32M LiDAR frames, 12.96M RGB, and 12.96M depth frames. This dataset also includes brightness, temperature, humidity, smoke, and airflow values covering all flight routes. 2) The platform supports the customization of simulators, UAVs, sensors, flight algorithms, formation modes, and ELH tasks. Through a visual control window, this platform allows users to collect customized datasets through one-click deployment online and to verify algorithms by closed-loop simulation, which can greatly alleviate the limitations of existing datasets on algorithm development. 3) U2UData+ also introduces an ELH task for wildlife conservation and provides 9 state-of-the-art swarm algorithms for benchmarking. All datasets, platforms, benchmarks, and video tutorials have been open-sourced and are available for public use. Our contributions can be summarized as follows:
@@
  • Dataset. We collect the first swarm UAV autonomous flight dataset for ELH tasks, with a size of over 3.62T.
  • Platform. We build the first scalable swarm UAV data online collection and algorithm closed-loop verification platform, which allows users to collect customized datasets and verify algorithms.
  • Benchmark. We introduce an ELH task for wildlife conservation and provide comprehensive benchmarks with 9 SOTA models.
Related Work
@@
Swarm UAV autonomous flight. Current low-altitude economy research mainly focuses on single-UAV autonomous flight and has matured core capabilities (5; 13; feng2024multi), including object detection, semantic segmentation, localization, obstacle avoidance, navigation, tracking, and stabilized flight control in controlled environments. However, they still suffer from many real-world challenges, for example: (1) Their perception remains fundamentally constrained by single-viewpoint occlusion and limited sensor range (zheng2025cross; li2023lot), severely reducing situational awareness in dynamic open environments. (2) Onboard computational resources restrict real-time decision-making for dynamic obstacle negotiation and executing ELH tasks (feng2025embodied; li2025miv; shen2025detach). (3) Operational robustness is inherently fragile (gao2025vagu; gao2025suvad; gao2025evolution), as hardware failures or unexpected obstacles often lead to task failure with no redundancy. Swarm UAV autonomous flight (12) can solve the inherent limitations of single-UAV through collaborative perception, localization, communication, navigation, tracking, and task re-allocation. Due to the lack of datasets, research on autonomous flight algorithms for swarm UAV has just begun.
@@
Swarm UAV simulators. Existing swarm UAV simulators include FightGear (S1), XPlan (S2), Jmavsim (S3), Gazebo (S4), AirSim (airsim), Rfly-Sim (S5), Isaac Sim (S6), and U2USim (u2usim). Swarm UAV simulators need to more realistically simulate dynamic physical characteristics (such as collision); sensors such as IMU, camera, GPS, LiDAR, temperature, humidity, and airflow due to their small size; and interaction with the ROS ecosystem. FightGear is not open source. XPlan and Jmavsim can only interact with ROS. Gazebo, AirSim, and RflySim can interact with ROS, simulate physical collision, and output visual sensor content. AirSim and RflySim can also implement weather control. However the information on these simulators is purely simulated, and the models trained on these simulators are difficult to run in the real world. Isaac Sim and U2USim add real environment data based on previous simulators. Isaac Sim can visually realize digital twins of the real world through GPU rendering, but it is difficult to provide modal information other than visual and LiDAR modalities. U2USim is the first real-world mapping swarm UAV simulator, taking Yunnan Province as the prototype, including 4 terrains, 7 weather conditions, and 8 sensor types. However, all parameters of U2USim are fixed: it only contains 3 types of animals, the number of animals is fixed, the intensity and range of weather are fixed, and the take-off point of the UAV is also fixed. If we want to test in another terrain, we need to fly to the target location for a long time before each test.
@@
Swarm UAV datasets. Public swarm UAV datasets have significantly accelerated progress in UAV flight technologies in recent years. As shown in Table 1, existing swarm UAV datasets include CoPerception-UAVs (17), CoPerception-UAVs+ (36), and U2UData (feng2024u2udata). CoPerception-UAVs (17) and CoPerception-UAVs+ (36) rely on open-source simulators like AirSim (airsim) and CARLA (CARLA), featuring limited terrain, weather, and sensor types. These datasets collect data at fixed altitudes and in consistent or fixed formation modes. In contrast to autonomous driving (liu2024survey), UAVs’ autonomous flight presents greater freedom, encounters more complex environments, and is more susceptible to the influence of temperature, humidity, and airflow due to its smaller size. Hence, there exists a notable domain gap between existing synthetic data and real-world data, potentially limiting the generalization of models trained. U2UData (feng2024u2udata) is the first large-scale cooperative perception dataset for swarm UAVs autonomous flight, which is collected by three UAVs flying autonomously in the U2USim (u2usim), covering a 9 km2 flight area, 4 terrains, 7 weather conditions, and 8 sensor types. U2UData manually selects 100 scenarios for each weather condition; U2UData collects 15 seconds of swarm UAV cooperative perception dataset for each scenario. Due to the emergence of U2UData, swarm UAV autonomous flight algorithms have begun to be studied. However, since U2UData only considers three UAVs tracking three animals, the length of each trajectory is only 15 seconds, and the dataset size and setting is fixed and cannot be expanded; only basic collaborative perception and tracking tasks can be designed. Complex ELH tasks (liu2025tcdformer; Wang_improving) for swarm UAVs in dynamic open environments cannot be explored.
@@
Figure 2: The scalable UAVs data online collection and algorithm closed-loop verification platform. Wind throughout the map.
U2UData+ Platform
U2UData+ bulids the first scalable swarm UAV data online collection and algorithm closed-loop verification platform, as shown in Figure 1, supporting the customization of simulators, UAVs, sensors, flight algorithms, formation modes, and ELH tasks. Through a visual control window, U2UData+ allows users to collect customized datasets through one-click deployment online and to verify algorithms by closed-loop simulation. We have built a video tutorial for this platform. For each scalable operation, users can complete it one by one according to the video tutorial.
Real-world mapping simulator. The platform is based on U2USim (u2usim), a real-world mapping swarm UAV simulator. The platform uses Unreal Engine (UE) 5.2111https://www.unrealengine.com/en-US/unreal-engine-5 to construct a scaled-down 3km*3km simulated environment map based on the map of Yunnan Province. The platform includes 4 types of terrain: mountains, hills, plains and basins. The elevation range is [56.6, 3000]m. Based on the vegetation and animal distribution in Yunnan, 58 types of original forest vegetation and 15 types of animal assets were constructed, and more than 15 superposition methods were used to combine vegetation assets, including epiphytic growth, diagonal staggered growth, and so on. Among them, the leaves of each plant will dynamically change with wind, snow, and other weather conditions. This platform can represent complex ecological relationships between animals. It accurately models interactions such as predator-prey, showing how these relationships influence collective movement patterns. The platform includes 7 weather conditions: sunny, rain, snow, sandstorm, wind, thunder, and fog at specific positions within the simulation environment. The platform uses the real meteorological data of Yunnan Province collected by the China Meteorological Center to map the simulation environment based on longitude and latitude. Among them, temperature and humidity are scalars, and missing values are filled by the moving average method (interval 5m). Wind speed and wind direction are first decomposed into scalars along longitude and latitude, then missing values are filled by sliding average, and finally constructed by vector synthesis.
@@
Table 2: A detailed comparison of the data size between U2UData+ with existing swarm UAV datasets.
@@
U2UData+ 12.96M 1920*1080 12.96M 4.32M 25.92M 12.96M 12.96M 12.96M 12.96M
Scalable ELH tasks. The platform provides four preset ELH tasks: wildlife conservation employs adaptive animal tracking algorithms using real-time behavior prediction across variable terrains and vegetation density. Logistics distribution dynamically reroutes paths around simulated urban obstacles and weather disruptions while maintaining payload integrity. Patrol security implements anomaly detection through continuous environmental scanning, adapting surveillance patterns to emergent threats in real-time. Disaster rescue prioritizes survivor identification in volatile conditions (collapsing structures, spreading fires) via multi-sensor fusion and probabilistic hazard mapping. Each task integrates specialized perception-action loops that respond to unpredictable environmental changes without predefined waypoints, such as sudden weather shifts or moving obstacles. New ELH tasks (e.g., precision agriculture) can be added by modifying the UE5.2 simulator source code.
Scalable flight algorithms and formation modes. The platform supports four swarm UAV autonomous flight algorithms for four ELH tasks: wildlife conservation, logistics distribution, patrol security, and disaster rescue. Those algorithms are built upon modularized components, including task planning, collaborative perception, localization, communication, navigation, tracking, and task re-allocation. Users can add or modify these algorithms via the open-source code of the visual control window, where modular code blocks allow drag-and-drop replacement or augmentation of existing logic. New autonomous flight algorithms for custom ELH tasks can be integrated by directly modifying the provided Python/ROS 2 interfaces in the code repository. The platform implements three distinct swarm formation modes: Discipline formation mode maintains strict geometric coordination (e.g., linear/radial arrays) for high-precision collaborative tasks, with real-time position correction compensating for environmental disturbances. Fixed formation mode enables individual UAVs to follow predefined paths, critical for infrastructure inspection or convoy protection scenarios. Autonomous formation mode supports dynamic reconfiguration where UAVs independently adapt spacing and topology using real-time perception data, ideal for complex environments like wildlife conservation or disaster rescue. Users can select any swarm formation modes via the visual control window for task-specific optimization.
Visual control window. The platform provides a visual control window, as shown in Figure 3. Users can collect datasets through a one-click deployment of the customized UAV starting point, weather, ELH tasks, swarm UAV autonomous flight algorithms, and swarm formation mode. Users can also verify swarm UAV autonomous flight algorithms by closed-loop simulation. For the platform basic capability test, users can first click the ”keyboardControl” button, and then control the UAV by clicking the following buttons: ”Take off”, ”Land”, ”Up”, ”Down”, ”Move Forward”, ”Move Backward”, ”Move Left”, and ”Move Right”. We’ve also implemented XBOX controller control. Connect your XBOX and open the simulator to directly control the UAV with the controller. It’s important to note that XBOX controller control and keyboard control are mutually exclusive.
U2UData+ Dataset
U2UData+ collects the first swarm UAV autonomous flight dataset for ELH tasks, with a size of over 3.62T. The dataset is collected by 15 UAVs in autonomous formation mode for the ELH task (wildlife conservation).
ELH tasks. U2UData+ only collects one ELH task: wildlife conservation, with a size of over 3.62T. Due to the huge amount of data, users can collect datasets for other ELH tasks on the U2UData+ platform on their own.
@@
Table 4: Data collection settings between U2UData+ with existing swarm UAV datasets. ET-Length: The length of each trajectory. TNT: The total length of trajectories.